A coin bank contains only dimes and quarters. The ToTal value of The coins in The bank is $3.70 . If the dimes were quarters and quarters were dimes, the total value of the coIns would be $2.95 . Find the number of quarters in the bank.
Answer by mananth(16949)

dimes = x numbers
quarters = y numbers
convert $ to cents
10x+25y=370
25x+10y=295
dimes x
Quarters y

10 x + 25 y = 370 .............1
25 x + 10 y = 295 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by -5
20 x 50 y = 740
-125 x + -50 y = -1475
Add the two equations
-105 x = -735.00
/ -105
x = 7
plug value of x in (1)
10 x + 25 y = 370
70 + 25 y = 370
25 y = 370 -70
25 y = 300
y = 12
dimes = 7
Quarters= 12
m.ananth@hotmail.ca